write an equation for the line containing the points (2, 2) and (3, -3)
Answer by Cromlix(4381)

Hi,
Slope formula: y2 - y1/x2 - x1
= -3 -2/3 - 2 = -5/1 = -5
Using formula: y - b = m(x - a) m = -5 and (a,b) = (2, 2)
y - 2 = -5(x - 2)
y - 2 = -5x + 10
y = -5x + 12.
Hope this helps :-)